OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-8, -11) and is perpendicular to -2x – 7y = -14 Please show all of your work

OpenStudy (anonymous):

we already have a point the line goes through so all we need is slope so we can put it in slope intercept form... since we know the line must be perpendicular to the line -2x - 7y = -14, the slope we want is the negative reciprocal of the slope of this line... -2x-7y=-14 -7y = -14 + 2x y = 2 - (2/7)x since the slope is -2/7, the slope we want is the negative reciprocal of this.. so the slope we want is 7/2. so now we need to find b in: y = mx + b plug in the given point: -11 = (7/2)(-8) + b -11 = -28 + b 17 = b so your equation is: y = (7/2)x + 17

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw...:)